An electron in an atom has principal quantum numbern 5. (a) What are the possible values of ? (Give the answer as a comma separated list.) (b) What are the possible combinations of t and m? (Do this on paper. Your instructor may ask you to turn in this work.) (c) Using the fact that there are two quantum states for each combination of and me because of electron spin, find the total number of electrons states for n 5. eBook
Answer #1

Poseuble values 1-1 tn-1 to o X= 4, 3, 2, 1, o nalleMs do J-千 ,m,--4,-3,-2,-1 , o , 1,2, 3,4 1-3, Yng=-3,-2,-1,0, 1, 2, 3 Lc)